# Topology-Driven Quantum Architecture Search Framework

### v1

@@ Abstract
The limitations of Noisy Intermediate-Scale Quantum (NISQ) devices have spurred the development of Variational Quantum Algorithms (VQAs), which hold the potential to achieve quantum advantage for specific tasks. Quantum Architecture Search (QAS) algorithms are pivotal in automating the design of high-performance Parameterized Quantum Circuits (PQCs) for VQAs. However, existing QAS algorithms grapple with excessively large search spaces, resulting in significant computational complexity when searching large-scale quantum circuits. To address this challenge, we propose a Topology-Driven Quantum Architecture Search (TD-QAS) framework. Our framework initially employs QAS to identify optimal circuit topologies and subsequently utilizes an efficient QAS to fine-tune gate types. In the fine-tuning phase, the QAS inherits parameters from the topology search phase, thereby eliminating the need to train from scratch. By decoupling the extensive search space into topology and gate-type components, TD-QAS avoids exploring gate configurations within low-performance topologies, significantly reducing computational complexity. Numerical simulations across various tasks, under both noiseless and noisy scenario, demonstrate that the TD-QAS framework enhances the efficiency of QAS algorithms by enabling them to identify high-performance quantum circuits with lower computational complexity. These results suggest that TD-QAS will deepen our understanding of VQA and have broad applicability in future QAS algorithms.
@@ 1 Introduction
Quantum computing exhibits promising potential in addressing complex problems such as combinatorial optimization [1, 2], factoring [3], linear system solving [4], chemical simulations [5], and machine learning [6, 7, 8, 9, 10, 11, 12]. However, Noisy Intermediate-Scale Quantum (NISQ) devices [13], representing the cutting edge of near-term quantum technology, encounter significant challenges, including noise, a limited number of qubits, and restricted qubit connectivity. These limitations impede the widespread application of quantum computing [14]. To mitigate these issues, Variational Quantum Algorithms (VQAs) have been developed to operate within the constraints of NISQ devices [15, 16, 17, 18], offering the potential to achieve quantum advantage for specific tasks.
@@
A critical factor in the effectiveness of VQAs is the design of Parameterized Quantum Circuits (PQCs) [19, 20]. Quantum Architecture Search (QAS) is a technique for automating the search for high-performing PQCs [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. QAS aim to explore the optimal quantum circuit in a vast search space encompassing all possible circuit configurations. Consequently, QAS algorithms face the challenge of excessive computational complexity due to the need to explore an exponentially large search space, particularly when targeting large-scale quantum circuits [31]. Therefore, developing methods to efficiently reduce the search space is essential for enhancing the efficiency and scalability of QAS.
@@
Quite a few methods have already been proposed to reduce the search space for QAS algorithms. A common strategy involves imposing rigid constraints on PQC design [32], such as restricting quantum gates to operate only on fixed qubits rather than arbitrary qubits. However, these methods indiscriminately discard a large number of potential PQCs without any targeted strategy, which may inadvertently eliminate optimal circuit. Another approach focuses on identifying components of quantum circuits and constructing larger circuits by repeating these components [27, 33]. This method is particularly suitable for managing large search spaces, although it may result in circuits with greater depth. Additionally, a preprocessing method has been proposed for QAS to eliminate low-performance quantum circuits [34]. While this method significantly improve the search efficiency of QAS, they introduce additional computational complexity.
@@
In this paper, we propose a novel framework to significantly reduce the search space in QAS. Our approach is inspired by Neural Architecture Search (NAS), which aims to automatically discover high-performance neural network architectures for a given task [36]. It is well established that the performance of Convolutional Neural Networks (CNNs) identified through NAS remains relatively robust to the random replacement of specific operations [37]. Building on this observation, we consider whether a similar phenomenon could be observed in VQA, where topology plays a more crucial role than specific gate types in determining performance. This motivates us to investigate a method of decoupling the QAS process into two independent components, i.e., topology and gate types. The decoupling method enables QAS to avoid exploring gate-type configurations associated with low-performance topologies.
@@
Through extensive numerical simulations, detailed in Section 4.1, we validate that topology indeed plays a dominant role over gate types in the search for high-performance quantum circuits. This finding suggests that by fixing the gate types and concentrating solely on topology, QAS can effectively identify high-performance quantum circuits. Furthermore, we observe that during the gate-type optimization phase, the trainable parameters of QAS can be inherited from the topology search phase. This inheritance eliminates the need for training from scratch, thereby accelerating convergence. Consequently, we propose the Topology-Driven Quantum Architecture Search (TD-QAS) framework, which prioritizes an independent search for high-performance topologies followed by an efficient fine-tuning of gate types. The TD-QAS framework effectively decouples the original search space into two distinct spaces, i.e., topology and gate types, thereby significantly reducing the overall search space. We conducted numerical simulations across a variety of tasks in both noiseless and noisy scenarios. These results demonstrate that the proposed TD-QAS framework enables QAS algorithms to identify quantum circuits with comparable performance while maintaining significantly low computational complexity, compared by original QAS algorithms that searching both topology and gate types. This study not only deepens our understanding of VQAs but also significantly extends the applicability of QAS to more complex tasks, paving the way for more efficient and scalable approaches to designing quantum circuits.
@@ 2.1 Quantum Architecture Search
QAS is a promising technique for automating the search of high-performing PQCs, capable of integrating specific target tasks or additional constraints during the search process [30]. It navigates the vast search space of potential quantum circuits by employing a search strategy to generate candidate circuits, followed by an evaluation strategy to assess their performance. The feedback from these performance is then used to optimize the search model’s trainable parameters. However, QAS faces significant challenges related to high computational complexity, prompting the development of various approaches to mitigate these expenses.
Efforts to reduce the computational complexity of QAS have concentrated on three key aspects: search strategy, evaluation strategy, and search space optimization. Search Strategy: The most basic search strategy is a random search method, which often demands an impractically large number of iterations to locate an effective circuit, leading to excessive computational complexity. To address this challenge, more sophisticated search models have been introduced to enhance search efficiency or reduce training costs. These advanced models include evolutionary algorithms, reinforcement learning [21], differentiable algorithms[24], Bayesian optimization methods[27], and so on. Evaluation Strategy: While these methods are effective in improving search efficiency, they necessitate substantial feedback for parameter training, resulting in the evaluation of an excessive number of quantum circuits. To mitigate the high computational burden associated with evaluating numerous quantum circuits, more efficient evaluation techniques have been developed. These methods aim to accelerate the evaluation process while striving to maintain evaluation accuracy. Notable approaches include shared parameter evaluation (one-shot techniques) [23], performance predictor-based evaluations [25], and training-free predictors [28]. Search Space Optimization: The extensive search space in QAS introduces additional challenges, not only in terms of the search model’s ability to efficiently learn from the exploration process but also in accurately evaluating the performance of quantum circuits. To overcome these obstacles, various techniques have been proposed to reduce the search space. These include imposing strict constraints on quantum gate operations [31], searching for small-scale circuit components to build larger circuits [27, 32], and employing search space pruning method [33]. By narrowing down the search space, these techniques enhance the efficiency and scalability of QAS, enabling it to focus on more promising regions of the search space.
In this work, we implement the TD-QAS framework using two representative QAS algorithms, i.e., QuantumSupernet [23] and DQAS [24], to validate the improvements brought by our framework over traditional QAS algorithms. The following subsections provide detailed introductions to these two algorithms.
@@ 2.1.1 QuantumSupernet
QuantumSupernet, proposed by Du et al. [23], is an advanced QAS algorithm designed to optimize the search process for high-performing quantum circuits. We introduce the QuantumSupernet by focusing on three QAS’s key aspects. QuantumSupernet employs a supernet to define the search space, which is composed of multiple layers containing single-qubit and two-qubit placeholders that act on all qubits. Each placeholder can be filled with specific gates from a native gate set, which includes predefined selectable quantum gates. This structured approach allows for comprehensive exploration within a manageable framework. The process of assigning specific quantum gates to all placeholders is referred to as sub-sampling. QuantumSupernet utilizes two distinct sub-sampling strategies: random sampling and a genetic algorithm-based search strategy. QuantumSupernet employs a shared parameter-based evaluation strategy to assess the performance of sampled sub-circuits. This method significantly accelerates the QAS process by allowing for the rapid evaluation of multiple sub-circuits without training parameters from scratch.
The QuantumSupernet workflow consists of four main phases. The first is the initialization phase, which defines the supernet structure and native gate set. The second phase is the shared parameter training phase, QuantumSupernet iteratively samples sub-circuits and assesses their performance using corresponding shared parameters, updating the shared parameter pool accordingly. Next is high-perforamance ansatzs search phase, A search strategy is employed to sample multiple sub-circuits, and their predicted performance is evaluated based on the shared parameter pool. This process outputs the sub-circuit with the highest predicted performance. The last is the output phase, the selected sub-circuit is trained from scratch to evaluate its actual performance, providing a final assessment of its efficacy.
@@ 2.1.2 DQAS
Differentiable Quantum Architecture Search (DQAS), proposed by Zhang et al., is the first to introduce a differentiable search model to accelerate the QAS algorithm [24]. We introduce the DQAS algorithm by focusing on three key aspects. Unlike QuantumSupernet, DQAS allows quantum gates to act on arbitrary qubits rather than strictly constraining their arrangement. DQAS employs a differentiable search strategy to sample multiple quantum gate operations, which are then used to construct complete quantum circuits. Specifically, DQAS utilizes a probability model, denoted as $P(\alpha)$, to guide the search process. This model transforms the discrete task of searching for quantum circuit architectures into a continuous optimization problem by adjusting the parameters of the probability model through gradient descent techniques. Similar to QuantumSupernet, DQAS adopts a shared parameter-based evaluation strategy to assess the performance of sampled circuits. This approach allows for the efficient evaluation of multiple circuit configurations without the need to train each circuit’s parameters from scratch, thereby reducing computational overhead.
The DQAS workflow consists of four main phases. In the initialization phase, defining the native gate set and the number of gates in the quantum circuit, setting the parameters for the search space. DQAS iteratively samples quantum circuits using the probability model and assesses their performance using shared parameters. This sampling process explores the search space by generating diverse circuit based on the current probability distribution. In the training phase, the performance of the sampled circuits is used to update both the shared parameter pool and the probability model through gradient descent. This phase iteratively refines the search strategy by adjusting the probability model to favor higher-performing circuit configurations. Last, in the output phase, Once the termination condition is met, the final circuit is trained from scratch to evaluate its actual performance. This trained circuit serves as the final output of the algorithm.
@@ 2.2 Numerical Simulation Tasks
VQAs play a crucial role in quantum computing, offering a wide range of important applications. In QAS, three representative tasks are commonly used as benchmarks to validate the effectiveness of QAS algorithms: ground state energy estimation for molecular systems, solving the MaxCut problem, and classification tasks.
@@ 2.2.1 Variational Quantum Eigensolver
Calculating the ground state and the lowest energy of Hamiltonian is a fundamental problem in physics research [38, 39, 40]. In quantum systems, the Hamiltonian’s dimensionality increases exponentially with the size of system, making it extremely difficult to address the problem. Therefore, researchers have employed the Variational Quantum Eigensolver (VQE), which offers exponential speedup, to address this issue [41, 42, 43, 44]. The key to solving this problem is to prepare the ground state $|\varphi_{\text{0}}\rangle$ of the Hamiltonian $H$ and compute the lowest energy $E_{\text{0}}$, as expressed in (1):
$E_{0}=\frac{\langle\varphi_{0}^{*}|H|\varphi_{0}\rangle}{\langle\varphi_{0}^{*}|\varphi_{0}\rangle}$ (1)
Since the $|\varphi_{0}\rangle$ is difficult to prepare, VQE approximates it by constructing a states $|\varphi^{n}\rangle$ by a PQC $U(\theta)$, as shown in (2):
$|\varphi^{n}\rangle=U(\theta)|0^{n}\rangle$ (2)
Here, $|0^{n}\rangle$ represents the all-zero quantum state, which often serves as the initial state is and processed by the $U(\theta)$. The goal of VQE is to iteratively adjust the parameters $\theta$ so that $|\varphi^{n}\rangle$ gradually approaches $|\varphi_{0}\rangle$. Therefore, calculating the ground state and the lowest energy can be transformed into finding a suitable $\theta$ that minimizes (3):
$\theta^{*}=\arg\min_{\theta}L(\theta)=\langle 0^{n}|U^{\dagger}(\theta)HU(\theta)|0^{n}\rangle$ (3)
In conclusion, The process of VQE can be summarized as first preparing the quantum state $|\varphi^{n}\rangle$ through $U(\theta)$, then measuring the energy $E$ of the state $|\varphi^{n}\rangle$, last update $\theta$ until the minimum $E$ is obtained.
@@ 2.2.2 Variational Quantum Algorithm for the MaxCut Problem
The MaxCut problem is a classic combinatorial optimization problem. Given a graph $G=(V,E)$, the objective is to partition its nodes into two distinct sets to maximize the number of edges connecting these sets. The objective function of MaxCut can then be written as:
$C(z)=\frac{1}{2}\sum_{(i,j)\in E}\left(1-z_{i}z_{j}\right)$ (4)
where $z_{i},z_{j}\in\{+1,-1\}$ and $z_{i}$ represents the subset to which vertex $i$ belongs. The Quantum Approximate Optimization Algorithm (QAOA) is commonly applied to address the MaxCut problem by utilizing alternating optimization to approximate the optimal solution, offering the potential for quantum speedup [1]. In the context of QAOA, the problem Hamiltonian $H_{c}$ that encodes the MaxCut objective can be expressed as:
$H_{C}=\frac{1}{2}\sum_{(i,j)\in E}\left(I-Z_{i}Z_{j}\right)$ (5)
where $Z_{i}$ and $Z_{j}$ are Pauli-Z operators acting on qubits corresponding to the nodes $i$ and $j$ of edge $(i,j)\in G$. These operators are used to encode the connectivity of the graph in the quantum circuit. The PQC of QAOA is driven by the problem. By adjusting PQC’s parameters, QAOA can progressively approximate the optimal solution. In this paper, we utilize a VQA searched by the QAS algorithm to replace ansatz of QAOA for solving the MaxCut problem. The solving process is similar to VQE, except for the Hamiltonian $H$.

### v2

@@ Abstract
The limitations of Noisy Intermediate-Scale Quantum (NISQ) devices have motivated the development of Variational Quantum Algorithms (VQAs), which are designed to potentially achieve quantum advantage for specific tasks. Quantum Architecture Search (QAS) algorithms play a critical role in automating the design of high-performance Parameterized Quantum Circuits (PQCs) for VQAs. However, existing QAS approaches struggle with large search spaces, leading to substantial computational overhead when optimizing large-scale quantum circuits. Extensive empirical analysis reveals that circuit topology has a greater impact on quantum circuit performance than gate types. Based on this insight, we propose the Topology-Driven Quantum Architecture Search (TD-QAS) framework, which first identifies optimal circuit topologies and then fine-tunes the gate types. In the fine-tuning phase, the QAS inherits parameters from the topology search phase, eliminating the need for training from scratch. By decoupling the large search space into separate topology and gate-type components, TD-QAS avoids exploring gate configurations within low-performance topologies, thereby significantly reducing computational complexity. Numerical simulations across various tasks, under both noiseless and noisy conditions, validate the effectiveness of the TD-QAS framework. This framework advances standard QAS algorithms by enabling the identification of high-performance quantum circuits while minimizing computational demands. These findings indicate that TD-QAS deepens our understanding of VQAs and offers broad potential for the development of future QAS algorithms.
@@ 1 Introduction
Quantum computing exhibits promising potential in addressing complex problems, such as combinatorial optimization [1, 2], factoring [3], linear system solving [4], chemical simulations [5], and machine learning [6, 7, 8, 9, 10, 11, 12]. However, Noisy Intermediate-Scale Quantum (NISQ) devices [13], which represent the forefront of near-term quantum technology, face substantial challenges, such as noise, limited qubit count, and restricted qubit connectivity. These limitations impede the widespread application of quantum computing [14]. Therefore, Variational Quantum Algorithms (VQAs) have been developed to operate within the constraints of NISQ devices [15, 16, 17, 18], offering the potential to achieve quantum advantage for specific tasks.
@@
The effectiveness of Variational Quantum Algorithms (VQAs) highly depends on the design of Parameterized Quantum Circuits (PQCs) [19, 20]. Quantum Architecture Search (QAS) is a technique employed to automate the discovery of high-performance PQCs [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. QAS aims to identify the optimal quantum circuit within a vast search space that encompasses all possible circuit configurations. Consequently, QAS algorithms encounter significant computational challenges due to the exponential growth of the search space, particularly when scaling to large quantum circuits [32]. Therefore, developing strategies to efficiently reduce the search space is crucial for improving the efficiency and scalability of QAS.
@@
Several methods have been proposed to reduce the search space for QAS algorithms. A common approach involves imposing strict constraints on the design of PQCs [33], such as limiting quantum gates to act on fixed qubits rather than arbitrary ones. However, these methods indiscriminately discard numerous potential PQCs without a targeted approach, which may inadvertently exclude optimal circuit configurations. Another approach focuses on identifying modular components of quantum circuits and constructing larger circuits by repeating these components [27, 34]. This method effectively manages large search spaces but may results in circuits with increased depth. Additionally, a preprocessing technique has been proposed to filter out low-performance quantum circuits in QAS [35]. While this approach significantly improves the search efficiency of QAS, it introduces additional computational complexity.
@@
In this paper, we propose a novel framework to significantly reduce the search space in QAS. Our approach is inspired by Neural Architecture Search (NAS), which aims to automatically identify high-performance neural network architectures for a specific task [37]. It has been established that the performance of Convolutional Neural Networks (CNNs) identified through NAS remains relatively robust to the random replacement of specific operations, such as convolutional layers [38]. Building on this observation, we explore whether a similar phenomenon occurs in VQA, where the circuit topology plays a more critical role than specific gate types in determining performance. This motivates us to investigate a method for decoupling the QAS process into two independent components: topology and gate types. This decoupling strategy allows QAS to avoid exploring gate-type configurations associated with low-performance topologies.
@@
Through extensive numerical simulations, we validate that topology indeed plays a more dominant role than gate types in the search for high-performance quantum circuits. Therefore, we propose the Topology-Driven Quantum Architecture Search (TD-QAS) framework, which prioritizes an independent search for high-performance topologies, followed by an efficient fine-tuning of gate types. The TD-QAS framework effectively decouples the original search space into two distinct spaces —topology and gate types —thereby significantly reducing the overall search space. Furthermore, we observe that during the gate-type optimization phase, QAS can inherit trainable parameters from the topology search phase. This inheritance method eliminates the need to train the gate-type optimization phase from scratch, accelerating convergence. We performed numerical simulations on various tasks in both noiseless and noisy scenarios. The results demonstrate that the proposed TD-QAS framework successfully identifies high-performance quantum circuits while maintaining significantly lower computational complexity compared to original QAS algorithms, which search both topology and gate types. This study enhances our understanding of VQAs as well as broadens the applicability of QAS to more complex tasks, paving the way for more efficient and scalable approaches to quantum circuit design.
@@ 2.1 Quantum architecture search
QAS is a promising technique for automating the search for high-performing PQCs, enabling integration of specific target tasks or additional constraints during the search process [30]. It explores the expansive search space of potential quantum circuits by employing a search strategy to generate candidate circuits, followed by an evaluation strategy to assess their performance. The feedback from these performance evaluations is used to optimize the trainable parameters of the search model. However, QAS faces significant challenges related to high computational complexity, which has spurred the development of various strategies to mitigate these costs.
Efforts to reduce the computational complexity of QAS have focused on three key aspects: search strategy, evaluation strategy, and search space optimization. The most basic search strategy is random search strategy, which often requires an impractically numerous iterations to identify an effective quantum circuit, causing excessive computational cost. Therefore, advanced search models have been introduced to enhance search efficiency or reduce training costs. These models include evolutionary algorithms, reinforcement learning [21], differentiable algorithms [24], and Bayesian optimization methods [27], among others.
Although advanced search strategies enhance search efficiency, they require substantial feedback for parameter training, leading to the evaluation of numerous quantum circuits; thus, more efficient evaluation techniques have been developed. These methods aim to accelerate the evaluation process while maintaining evaluation accuracy. Notable approaches include shared parameter evaluation (one-shot techniques) [23], performance predictor-based evaluations [25, 28], and training-free predictors [22].
The extensive search space in QAS presents challenges in efficiently learning from exploration and accurately evaluating quantum circuit performance. Therefore, various techniques have been proposed to reduce the search space. These include imposing strict constraints on quantum gate operations [32], searching for small-scale circuit components to construct larger circuits [27, 33], and employing search space pruning methods [34]. By narrowing the search space, these techniques improves QAS efficiency and scalability, enabling it to focus on more promising regions.
In this work, we implement the TD-QAS framework based on two representative QAS algorithms (QuantumSupernet [23] and DQAS [24]), to validate the improvements introduced by our approach over traditional QAS methods. The following subsections provide concise overview of the two baseline algorithms, while Section 4.3 describes how TD-QAS can be seamlessly integrated with them without altering their core mechanisms.
@@ 2.1.1 QuantumSupernet
QuantumSupernet, proposed by Du et al. [23], is a representative QAS algorithm designed to efficiently explore large quantum circuit spaces using a one-shot neural architecture search paradigm. The core concept of QuantumSupernet is to construct a parameter-sharing supernet that encodes numerous candidate circuits within a unified structure. The supernet consists of multiple layers, each containing a set of single-qubit and two-qubit gates from a predefined native gate set. This layered representation offers a flexible yet efficient search space for quantum circuit design. To explore this space, QuantumSupernet employs a sub-sampling mechanism to extract candidate sub-circuits. The method supports both uniform random sampling and genetic algorithm-based sampling strategies. The performance of each sub-circuit is estimated through a shared-parameter evaluation mechanism, eliminating the need to train every candidate circuit from scratch. The shared parameters are jointly optimized within the supernet, enabling efficient reuse and rapid evaluation of multiple sub-circuits.
The overall workflow of QuantumSupernet comprises four stages: (1) initialization phase, in which the supernet architecture and native gate set are defined; (2) shared-parameter training phase, where sub-circuits are iteratively sampled and evaluated using the shared parameters, which are updated based on performance feedback; (3) architecture search phase, during which multiple sub-circuits are generated using either random or genetic sampling strategies, which are evaluated using shared parameters; (4) evaluation phase, where the best-performing sub-circuit is retrained from scratch to obtain its actual performance. This streamlined process enables QuantumSupernet to significantly reduce the computational cost of QAS while maintaining competitive accuracy across a variety of quantum tasks.
@@ 2.1.2 DQAS
Differentiable Quantum Architecture Search (DQAS), proposed by Zhang et al. [24], is a flexible framework that incorporates differentiability into quantum circuit architecture search, enabling gradient-based optimization of circuit structures. DQAS distinguishes itself from other QAS algorithms by formulating the search for circuit structures as a differentiable process, thereby significantly enhancing search efficiency.
DQAS models quantum circuit design as a bi-level optimization problem, jointly optimizing both the circuit structure and the associated gate parameters. The core concept is to relax the discrete search space of circuit architectures into a continuous domain using a parameterized probabilistic search strategy. The search space consists of a pool of candidate operations (quantum gates) and a circuit encoding mechanism that enables flexible assignment of these operations to the architecture. During each training iteration, DQAS samples a batch of circuit structures from the probabilistic model and assess their performance using a shared-parameter evaluation strategy. Gradients for both structure parameters (which define the probabilities of selecting specific gates at each layer) and shared parameters (which define the variational parameters within each gate) are computed and updated via automatic differentiation. This approach allows efficient, end-to-end learning of both the circuit layout and gate parameters. The final architecture is determined by selecting the most probable circuit structure based on the learned probabilistic model. This structure is retrained from scratch to obtain its true performance. The overall DQAS workflow consists of five stages: initializing the gate pool, sampling circuit architectures, updating structure and gate parameters via gradient descent, selecting the optimal architecture, and retraining it from scratch. Further details are available in the original work [24].
@@ 2.2 Numerical simulation tasks
VQAs are fundamental to quantum computing, offering a wide range of important applications. In QAS, three representative benchmark tasks commonly used to validate algorithm effectiveness include ground state energy estimation for molecular systems, solving the MaxCut problem, and performing classification.
@@ 2.2.1 Variational quantum eigensolver
Calculating the ground state and the lowest energy of Hamiltonian is a fundamental problem in physics research [39, 40, 41]. In quantum systems, the Hamiltonian’s dimensionality increases exponentially with system size, posing significant computational challenges. Therefore, researchers have employed the Variational Quantum Eigensolver (VQE), which offers exponential speedup, to address this issue [42, 43, 44, 45]. The key to solving this problem is to prepare the ground state $|\varphi_{\text{0}}\rangle$ of the Hamiltonian $H$ and compute the lowest energy $E_{\text{0}}$, as expressed in (1):
$E_{0}=\frac{\langle\varphi_{0}^{*}|H|\varphi_{0}\rangle}{\langle\varphi_{0}^{*}|\varphi_{0}\rangle},$ (1)
As the $|\varphi_{0}\rangle$ is difficult to prepare, VQE approximates it by constructing a states $|\varphi^{n}\rangle$ by a PQC $U(\theta)$, as shown in (2):
$|\varphi^{n}\rangle=U(\theta)|0^{n}\rangle,$ (2)
Here, $|0^{n}\rangle$ represents the all-zero quantum state, which serves as the initial state and processed by the $U(\theta)$. The goal of VQE is to iteratively adjust the parameters $\theta$ such that $|\varphi^{n}\rangle$ gradually approaches $|\varphi_{0}\rangle$. Therefore, calculating the ground state and the lowest energy can be transformed into finding a suitable $\theta$ that minimizes (3):
$\theta^{*}=\arg\min_{\theta}L(\theta)=\langle 0^{n}|U^{\dagger}(\theta)HU(\theta)|0^{n}\rangle,$ (3)
The process of VQE can be summarized as first preparing the quantum state $|\varphi^{n}\rangle$ through $U(\theta)$, then measuring the energy $E$ of the state $|\varphi^{n}\rangle$, and finally updating $\theta$ until the minimum $E$ is obtained.
@@ 2.2.2 Variational quantum algorithm for the MaxCut problem
The MaxCut problem is a classic combinatorial optimization problem. Consider a graph $G=(V,E)$, the objective is to partition its nodes into two distinct sets to maximize the number of edges connecting these sets. The objective function of MaxCut can be written as:
$C(z)=\frac{1}{2}\sum_{(i,j)\in E}\left(1-z_{i}z_{j}\right),$ (4)
where $z_{i},z_{j}\in\{+1,-1\}$ and $z_{i}$ represents the subset to which vertex $i$ belongs. The Quantum Approximate Optimization Algorithm (QAOA) addresses the MaxCut problem by utilizing alternating optimization to approximate the optimal solution, offering the potential for quantum speedup [1]. Consequently, the problem Hamiltonian $H_{c}$ that encodes the MaxCut objective can be expressed as:
$H_{C}=\frac{1}{2}\sum_{(i,j)\in E}\left(I-Z_{i}Z_{j}\right),$ (5)
where $Z_{i}$ and $Z_{j}$ are Pauli-Z operators acting on qubits corresponding to the nodes $i$ and $j$ of edge $(i,j)\in G$. These operators are used to encode the connectivity of the graph in the quantum circuit. The PQC of QAOA is driven by the problem. By adjusting PQC’s parameters, QAOA can progressively approximate the optimal solution. In this study, we employed a VQA searched by the QAS algorithm to replace ansatz of QAOA for solving the MaxCut problem. The solving process is similar to VQE, except for the Hamiltonian $H$.